the ratio of weight on the moon to weight on earth is 1:6. how much would a 210-pound person weigh on the moon?

Answer by stanbon(75887)

the ratio of weight on the moon to weight on earth is 1:6. how much would a 210-pound person weigh on the moon?
-------
x/210 = 1/6
----
x = 210(1/6)
x = 35 lb
------------
Cheers,
Stan H.